OpenStudy (anonymous):

Anyone help me on this equation. 4x-5y=27 3x-9y=36

OpenStudy (anonymous):

Scale up so that you end up with: 12x-15y=81 12x-36y=144 Then factorise out your 12x so 81+15y=144+36y solve for y y=-3 sub back into formula for x hope this helped!

OpenStudy (nurali):

4x-5y=27-----(1) 3x-9y=36-----(2) Multiply euation '1' by 3 and equation '2' by 4 and subtract 12x-15y=81 12x-36y=144 ------------- 21y=-63 y=-3 put y in equation '2' 3x+27=36 3x=9 x=3